f(1)=3 , f’(1)=2 , then d/dx{log f(e+ 2x)} at x =0 is
  1. 2/3
  2. 3/2
  3. 2
  4. 0

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
on differentiation
 
f’(e^x + 2x) ( e^x + 2)/ f(e^x + 2x) 
 
Now put x = 0
 
f’(1) (3) / f(1) = 3* 2 /3 = 2